# Scalable piecewise smoothing with BART

### v3

@@ C.1 Ridge approximations anisotropic Hölder class functions
To approximate $f_{0}$ globally, we define a smooth partition of unity $\{\phi_{r}\}_{r=1}^{R}$ subordinate to the partition $\{\Psi_{r}\}$ such that:
  • (1)
    $\phi_{r}\in C^{\infty}([0,1]^{p})$
  • (2)
    $\sum_{r=1}^{R}\phi_{r}(\bm{x})=1,\ \forall\bm{x}\in[0,1]^{p}$
  • (3)
    $\phi_{r}:[0,1]^{p}\longrightarrow[0,1]$ is compactly supported on an extended version of $\Psi_{r}$.
The global approximant is constructed as
$f_{D}(\bm{x})=\sum_{r=1}^{R}\phi_{r}(\bm{x})f_{D}^{r}(\bm{x})$
Note that we define the smooth partition of unity is obtained by normalizing:
$\phi_{r}(\bm{x})=\frac{\tilde{\rho}_{r}(\bm{x})}{\sum_{s=1}^{R}\tilde{\rho}_{s}(\bm{x})}.$
where the smooth bump functions are defined in Chapter 2 of Lee, (2003):
$\tilde{\rho}_{r}(\bm{x})=\exp\left(-\frac{1}{1-\|\bm{x}-\bm{c}_{r}\|^{2}/\delta_{r}^{2}}\right),\quad\text{for }\|\bm{x}-\bm{c}_{r}\|<\delta_{r}.$
However, this construction does not guarantee that $f_{D}$ is a ridge function. In that regard, one can directly approximate $\phi_{r}$ or approximate an unnormalized bump function $\tilde{\rho}_{r}$ and then normalize. Either way, the key is that each $\phi_{r}$ can be realized by a finite sum of ridge units up to an arbitrarily small uniform error. To that end, we approximate each bump $\tilde{\rho}_{r}$ by a sum of ridge functions. Since $\tilde{\rho}_{r}\in C^{\infty}$, it is known for many analytic activation functions that we can approximate $\tilde{\rho}_{r}$ uniformly on $[0,1]^{p}$ to an accuracy $\delta$ using $D^{\prime}$ ridge units (Barron,, 1993; Pinkus,, 1997). That is,
$\|\tilde{\rho}_{r}-\tilde{\rho}^{\text{ridge}}\|_{\infty}\leq\delta_{r},\ \text{where}\ \tilde{\rho}^{\text{ridge}}(\bm{x})=\sum_{l=1}^{D^{\prime}}\beta_{l}^{(r)}\varphi(\bm{\omega}_{l}^{(r)}\cdot\bm{x}+b_{l}^{(r)})$
Choose $D^{\prime}$ such that $\delta_{r}\sim(D^{\prime})^{-\kappa}$ for some $\kappa>0$ depending on the smoothness of $\tilde{\rho}_{r}$.
Thus, we can form the approximate partition of unity $\{\widehat{\phi}_{r}\}$ by defining
$\widehat{\phi}_{r}(\bm{x})\;=\;\frac{\widetilde{\rho}_{r}^{\,\mathrm{ridge}}(\bm{x})}{\sum_{s=1}^{R}\widetilde{\rho}_{s}^{\,\mathrm{ridge}}(\bm{x})}.$
Provided $\sum_{s=1}^{R}\widetilde{\rho}_{s}^{\,\mathrm{ridge}}(\bm{x})\geq\gamma>0$ on $[0,1]^{p}$, ensuring positivity, and $\{\widehat{\phi}_{r}(\bm{x})\}$ is still a partition of unity.
Now, we write
$|\hat{\phi}_{r}(\bm{x})-\phi_{r}(\bm{x})|$ $=\left|\frac{\widetilde{\rho}_{r}^{\,\mathrm{ridge}}(\bm{x})}{\sum_{s}\widetilde{\rho}_{s}^{\,\mathrm{ridge}}(\bm{x})}-\frac{\tilde{\rho}_{r}(\bm{x})}{\sum_{s}\tilde{\rho}_{s}(\bm{x})}\right|$
$=\left|\frac{\widetilde{\rho}_{r}^{\,\mathrm{ridge}}(\bm{x})\sum_{s}\tilde{\rho}_{s}(\bm{x})-\tilde{\rho}_{r}(\bm{x})\sum_{s}\widetilde{\rho}_{s}^{\,\mathrm{ridge}}(\bm{x})}{(\sum_{s}\widetilde{\rho}_{s}^{\,\mathrm{ridge}}(\bm{x}))(\sum_{s}\tilde{\rho}_{s}(\bm{x}))}\right|$
Using triangle inequality on the numerator, since $\|\tilde{\rho}_{r}-\widetilde{\rho}_{r}^{\,\mathrm{ridge}}\|_{\infty}\leq\delta_{r}$ for each $r$, we get $|\tilde{\rho}_{r}(\bm{x})-\widetilde{\rho}_{r}^{\,\mathrm{ridge}}(\bm{x})|\leq\delta_{r}$ and similarly for sums over all $s$. Altogether, we obtain:
$\|\hat{\phi}_{r}-\phi_{r}\|_{\infty}\leq C\delta,$ (C3)
for some constant $C$ that depends on $\min_{s}\|\tilde{\rho}_{s}\|_{\infty}$ and $\delta=\max_{r}\delta_{r}$. Hence, each $\hat{\phi}_{r}$ can be made as close to $\phi_{r}$ as we wish by choosing $D^{\prime}$ large enough. Having approximated $\phi_{r}$ by $\widehat{\phi}_{r}$, we define the final global approximation as
$\widehat{f}(\bm{x})~{}\coloneqq~{}\sum_{r=1}^{R}\bigl{[}\widehat{\phi}_{r}(\bm{x})\bigr{]}\bigl{[}f_{D}^{r}(\bm{x})\bigr{]},$
where each $f_{D}^{r}$ approximates $f_{0}$ on subdomain $\Psi_{r}$. To see the total error, write:
$\|\,f_{0}-\widehat{f}\|_{\infty}~{}\leq~{}\bigl{\|}f_{0}(\bm{x})-\sum_{r}\phi_{r}(\bm{x})f_{D}^{r}(\bm{x})\bigr{\|}_{\infty}~{}+~{}\bigl{\|}\sum_{r}\phi_{r}(\bm{x})f_{D}^{r}(\bm{x})-\sum_{r}\widehat{\phi}_{r}(\bm{x})f_{D}^{r}(\bm{x})\bigr{\|}_{\infty}.$
We call these (I) and (II):
  1. (I)
    By definition of $\phi_{r}(\bm{x})$,
    $f_{0}(\bm{x})~{}-~{}\sum_{r}\phi_{r}(\bm{x})\,f_{D}^{r}(\bm{x})~{}=~{}\sum_{r}\phi_{r}(\bm{x})\bigl{[}f_{0}(\bm{x})~{}-~{}f_{D}^{r}(\bm{x})\bigr{]}.$
    Thus, taking absolute values and using the triangle inequality,
    $\bigl{\lvert}f_{0}(\bm{x})-\sum_{r}\phi_{r}(\bm{x})\,f_{D}^{r}(\bm{x})\bigr{\rvert}~{}\leq~{}\sum_{r}\phi_{r}(\bm{x})\,\bigl{\lvert}f_{0}(\bm{x})-f_{D}^{r}(\bm{x})\bigr{\rvert}.$
    From Lemma C1, each subdomain $\Psi_{r}$ admits
    $\sup_{\bm{x}\in\Psi_{r}}\bigl{\lvert}f_{0}(\bm{x})-f_{D}^{r}(\bm{x})\bigr{\rvert}~{}\leq~{}C_{r}\,D^{-\min_{j}\{\alpha_{r,j}\}/p}.$
    Hence, for any $\bm{x}\in\Psi_{r}$, we have $\bigl{\lvert}f_{0}(\bm{x})-f_{D}^{r}(\bm{x})\bigr{\rvert}\leq C_{r}\,D^{-\min_{j}\{\alpha_{r,j}\}/p}.$ Since $\phi_{r}(\bm{x})\leq 1$, we get
    $\sum_{r}\phi_{r}(\bm{x})\bigl{\lvert}f_{0}(\bm{x})-f_{D}^{r}(\bm{x})\bigr{\rvert}~{}\leq~{}\max_{1\leq r\leq R}\sup_{\bm{x}\in\Psi_{r}}\bigl{\lvert}f_{0}(\bm{x})-f_{D}^{r}(\bm{x})\bigr{\rvert}~{}\leq~{}\max_{r}C_{r}\,D^{-\min_{j}\{\alpha_{r,j}\}/p}.$
    Taking the supremum over $\bm{x}\in[0,1]^{p}$, we conclude
    $\sup_{\bm{x}\in[0,1]^{p}}\bigl{\lvert}f_{0}(\bm{x})-\sum_{r}\phi_{r}(\bm{x})\,f_{D}^{r}(\bm{x})\bigr{\rvert}~{}\leq~{}\max_{1\leq r\leq R}\{C_{r}\}\,D^{-\min_{r,j}\{\alpha_{r,j}\}/p}.$
    Letting $C=\max_{r}C_{r}R^{\min_{r,j}\alpha_{r,j}/p}$, we obtain
    $\bigl{\|}f_{0}(\bm{x})-\sum_{r}\phi_{r}(\bm{x})\,f_{D}^{r}(\bm{x})\bigr{\|}_{\infty}~{}\leq~{}C\,D^{\star\ -\min_{r,j}\alpha_{r,j}/p}.$
  2. (II)
    $\Bigl{\|}\,\sum_{r}\bigl{[}\phi_{r}(\bm{x})-\widehat{\phi}_{r}(\bm{x})\bigr{]}\,f_{D}^{r}(\bm{x})\Bigr{\|}_{\infty}~{}\leq~{}\sum_{r}\|\phi_{r}-\widehat{\phi}_{r}\|_{\infty}\;\sup_{\bm{x}}\bigl{|}\,f_{D}^{r}(\bm{x})\bigr{|}.$ Each $f_{D}^{r}$ is bounded by some constant depending on $\|f_{0}\|_{\infty},\lambda$. Then $\|\phi_{r}-\widehat{\phi}_{r}\|_{\infty}\leq C^{\prime}\delta$ from Equation C3 implies (II) is at most $O(\delta)$.
Hence,
$\|\,f-\widehat{f}\|_{\infty}~{}\leq~{}C\,D^{\star\ -\min_{r,j}\alpha_{r,j}/p}~{}+~{}O(\delta).$
By choosing $\delta\approx D^{\star\ -\min_{r,j}\alpha_{r,j}/p}$, the two terms become comparable, giving $\|\,f-\widehat{f}\|_{\infty}\approx O(D^{\star\ -\min_{r,j}\alpha_{r,j}/p})$. ∎
@@ C.2 Proof of Theorem 1
    The split-net $\mathcal{Z}$ is $(\chi_{0}^{*},c_{n})$-dense, where $c_{n}\lesssim n^{-\nicefrac{{(1+2a_{2})}}{{3a_{1}}}}(\log n)^{-\nicefrac{{1}}{{3a_{1}}}}$.
@@
    The split net is $(\Xi_{r}^{*},\bm{\alpha}_{r},L)$-regular for every $r=1,2,...,R$, with $2^{L}\asymp(\nicefrac{{n(\lambda p)^{2}}}{{(R\log n)}})^{\nicefrac{{p}}{{(2\bar{\alpha}+p}})}$.
@@
These assumptions ensure that the boundaries of the underlying box partition can be well detected by the binary tree partitioning rule. In particular, assumptions (A4) and (A6) ensures the split-net is dense enough to approximate the unknown partition’s boundaries. Assumption (A5) sets the anisotropic k-d tree depth $L$ so that the partition can achieve the correct approximation rate w.r.t. $\bm{\alpha}_{r}$. Assumption (A7) caps the maximum depth needed for any sub-box $\Omega_{r}^{\star}$ to $\lesssim\log n$, preventing excessively large trees. It is useful to state that throughout this section, we define the class of functions defined on any tree partition $\mathcal{T}$ as $\mathcal{F}_{\mathcal{T}}$.
@@
Assume (A1)–(A7), (P1)–(P4). Then for any $f_{0}\in\mathcal{H}^{\mathcal{A},p}_{\lambda}(\bm{\chi_{0}^{\star}})$, there exists an anisotropic k-d tree partition $\hat{\mathcal{T}}$ of depth $L$ such that $2^{L}\asymp(\nicefrac{{n(\lambda p)^{2}}}{{(R\log n)}})^{\nicefrac{{p}}{{(2\bar{\alpha}+p}})}$ and $\hat{f}_{0}\in\mathcal{F}_{\hat{\mathcal{T}}}$ such that
$\|f_{0}-\hat{f}_{0}\|_{\infty}\lesssim\epsilon_{n}\coloneqq(\lambda p)^{\nicefrac{{p}}{{(2\bar{\alpha}+p)}}}(\nicefrac{{(R\log n)}}{{n}})^{\nicefrac{{\bar{\alpha}}}{{(2\bar{\alpha}+p)}}}.$
@@
Fix $r$ and $k$. For any $\bm{x}\in\Omega_{rk}^{0}$, we define
$\bm{x}^{*}=\arg\min_{\bm{z}\in\text{cl}(\Omega_{rk}^{0}\cap\Xi_{r}^{*})}\|\bm{x}-\bm{z}\|_{1}.$
Observe the following string of inequalities:
$\lVert f_{0}(\bm{x})-\hat{f}_{0}(\bm{x})\rVert_{\infty}$ $=\lVert f_{0}(\bm{x})-f_{0}(\bm{x}^{*})+f_{0}(\bm{x}^{*})-\hat{f}_{0}(\bm{x})\rVert_{\infty}$
$\leq\lVert f_{0}(\bm{x})-f_{0}(\bm{x}^{*})\rVert_{\infty}+\lVert f_{0}(\bm{x})-\hat{f}_{0}(\bm{x})\rVert_{\infty}$
$\leq\underbrace{\lVert f_{0}(\bm{x})-f_{0}(\bm{x}^{*})\rVert_{\infty}}_{(i)}+\underbrace{\lVert f_{0}(\bm{x}^{*})-f_{0}(\bm{x})\rVert_{\infty}}_{(ii)}+\underbrace{\lVert f_{0}(\bm{x})-\hat{f}_{0}(\bm{x})\rVert_{\infty}}_{(iii)}$
Since we know that $f_{0}$ is $\alpha_{rj}$-Holder continuous in each box, this implies
$(i)\leq\lambda pc_{n}^{\min\alpha_{rj}}$
Since $\bm{x},\bm{x}^{*}\in\Omega_{rk}^{0}$, from the anisotropic tree construction, we have
$(ii)\leq\lambda p2^{-\nicefrac{{\bar{\alpha}L}}{{p}}}$
following the tree construction elaborated in the proof of Theorem 1 in Jeong and Ročková, (2023). Note that if $\left(l_{r1},...,l_{rp}\right)^{\top}$ be the counter vector returned by the anisotropic k-d tree construction algorithm, such that $L=\sum_{j=1}^{p}l_{rj}$, for $r=1,2,...,R$, then following the proof technique of Jeong and Ročková, (2023), we take $\tilde{l}_{rj}=\nicefrac{{L\bar{\alpha}}}{{p\alpha_{rj}}}$, which implies the equalities
$\alpha_{r1}\tilde{l}_{r1}=...=\alpha_{rp}\tilde{l}_{rp},\ \text{and}\ L=\sum_{j=1}^{p}\tilde{l}_{rj}$
Observing that $l_{rj}>\tilde{l}_{rj}-1$, we can claim that the above inequality holds since,
$(ii)\lesssim\lambda\sum_{j=1}^{p}2^{-\alpha_{rj}l_{rj}}\leq 2\lambda\sum_{j=1}^{p}2^{-\alpha_{rj}\tilde{l}_{rj}}\leq 2\lambda p2^{-\nicefrac{{\bar{\alpha}L}}{{p}}}$
Assumptions (A1) – (A4) imply that the Hölder function class can be well approximated by the class of ridge functions.
$(iii)=\lVert f_{0}(\bm{x})-\hat{f}_{0}(\bm{x})\rVert_{\infty}\leq CD^{\star\ -\min_{r,j}\alpha_{r,j}/p}$
where $C$ is a constant independent of $f_{0}$. Collating all the bounds, and observing that the upper bound for (iii) is asymptotically constant as $p\rightarrow\infty$, we have
$\|f_{0}-\hat{f}_{0}\|_{\infty}\lesssim\lambda p(c_{n}^{\min\alpha_{rj}}+2^{-\nicefrac{{\bar{\alpha}L}}{{p}}})\lesssim\epsilon_{n}$
Note that following the lines of Corollary 1 in Jeong and Ročková, (2023), if we assume (A2) and (A4), then $c_{n}^{\min_{r,j}\alpha_{r,j}}\lesssim\epsilon_{n}/\lambda p$ proves the final inequality. ∎

### v4

@@ C.1 Ridge approximations anisotropic Hölder class functions
The global ridge approximation proof proceeds in three main steps. First, we construct an ideal global approximant, $f_{\textnormal{ideal}}(\bm{x})$, by smoothly “gluing together” the local polynomial approximations from the proof of Lemma C1. We will bound the error between the true function $f_{0}$ and this ideal function. Second, we recognize that this ideal approximant is a single, globally smooth function. We can therefore apply universal approximation theorems to show that it can be well-approximated by our target structure: a single sum of ridge functions, which we call $\hat{f}(\bm{x})$. Finally, we combine the error terms and manage the complexity budget ($D$ vs. $D^{\star}$) to arrive at the final result.
1. Constructing and bounding the ideal approximant. Let the domain $\bm{\Psi}=[0,1]^{p}$ be partitioned into $R$ rectangular boxes $\{\Psi_{r}\}_{r=1}^{R}$. From the proof of Lemma C1, we know that for each subdomain $\Psi_{r}$, there exists a multivariate polynomial $P_{N,r}(\bm{x})$ of degree less than $N$ in each coordinate that approximates $f_{0}$ with the following error:
$\sup_{\bm{x}\in\Psi_{r}}|f_{0}(\bm{x})-P_{N,r}(\bm{x})|\leq C_{r}N^{-\min_{j}\alpha_{r,j}}$
To create a single global approximant from these local polynomials, we introduce a smooth partition of unity. This is a set of functions $\{\phi_{r}(\bm{x})\}_{r=1}^{R}$ such that: (i) Each $\phi_{r}$ is infinitely differentiable ($C^{\infty}$), (ii) $\phi_{r}(\bm{x})\geq 0$ for all $\bm{x}$, and $\phi_{r}(\bm{x})$ is zero outside a small neighborhood of $\Psi_{r}$. Additionally, (iii) $\sum_{r=1}^{R}\phi_{r}(\bm{x})=1$ for all $\bm{x}\in\bm{\Psi}$. This partition-of-unity technique is not uncommon in approximation theory, for instance, Radial-basis-function (RBF) solvers often build local RBF interpolants on overlapping patches, then blend them with a smooth partition-of-unity to get a global interpolant (Bernal et al.,, 2024). Using this partition, we define our ideal global approximant as:
$f_{\text{ideal}}(\bm{x})=\sum_{r=1}^{R}\phi_{r}(\bm{x})P_{N,r}(\bm{x})$
We now bound the error of this ideal approximant. Since $\sum_{r}\phi_{r}(\bm{x})=1$, we can write $f_{0}(\bm{x})=\sum_{r}\phi_{r}(\bm{x})f_{0}(\bm{x})$.
$|f_{0}(\bm{x})-f_{\text{ideal}}(\bm{x})|$ $=\left|\sum_{r=1}^{R}\phi_{r}(\bm{x})f_{0}(\bm{x})-\sum_{r=1}^{R}\phi_{r}(\bm{x})P_{N,r}(\bm{x})\right|$
$=\left|\sum_{r=1}^{R}\phi_{r}(\bm{x})\left[f_{0}(\bm{x})-P_{N,r}(\bm{x})\right]\right|$
$\leq\sum_{r=1}^{R}\phi_{r}(\bm{x})\left|f_{0}(\bm{x})-P_{N,r}(\bm{x})\right|$
Taking the supremum over all $\bm{x}$:
$\lVert f_{0}-f_{\text{ideal}}\rVert_{\infty}$ $\leq\sup_{\bm{x}}\sum_{r=1}^{R}\phi_{r}(\bm{x})\max_{s=1,\dots,R}\left(\sup_{\bm{y}\in\Psi_{s}}|f_{0}(\bm{y})-P_{N,s}(\bm{y})|\right)$
$=\left(\sup_{\bm{x}}\sum_{r=1}^{R}\phi_{r}(\bm{x})\right)\left(\max_{s=1,\dots,R}\lVert f_{0}-P_{N,s}\rVert_{\infty,\Psi_{s}}\right)$
$\leq\max_{r=1,\dots,R}\left(C_{r}N^{-\min_{j}\alpha_{r,j}}\right)$
$\leq C^{\prime}N^{-\min_{r,j}\alpha_{r,j}}$
2. Approximating the ideal function with ridge functions. The function $f_{\text{ideal}}(\bm{x})$ is a sum of products of $C^{\infty}$ functions ($\phi_{r}$) and polynomials ($P_{N,r}$), so it is itself a single, globally smooth ($C^{\infty}$) function. Crucially, it does not have the target structure of a simple sum of ridge functions.
Our goal is to find a function $\hat{f}(\bm{x})=\sum_{d=1}^{D^{\star}}\beta_{d}\varphi(\bm{\omega}_{d}^{\top}\bm{x}+b_{d})$ that approximates $f_{\text{ideal}}(\bm{x})$. Again observing that $f_{\text{ideal}}$ is continuous and $\hat{f}(\bm{x})$ is a single hidden layer of a neural network, Leshno et al., (1993, Theorem 1) guarantees that there exists an $\hat{f}(\bm{x})$ such that:
$\lVert f_{\text{ideal}}-\hat{f}\rVert_{\infty}\leq\epsilon^{\prime}$
The number of ridge units, $D^{\star}$, required to achieve this error $\epsilon^{\prime}$ depends on the complexity of the target function $f_{\text{ideal}}$, which is determined by the $R$ polynomials, each having $O(N^{p})$ coefficients. The total complexity of $f_{\text{ideal}}$ is roughly $O(R\cdot N^{p})$. We set our total ridge function budget $D^{\star}$ to be proportional to this complexity:
$D^{\star}\asymp R\cdot N^{p}\implies N\asymp(D^{\star}/R)^{1/p}$
With this budget, it is standard from approximation theory that we can choose $\hat{f}$ such that the approximation error is of the same order as the error from Step 1. We set:
$\lVert f_{\text{ideal}}-\hat{f}\rVert_{\infty}\leq C^{\prime\prime}(D^{\star}/R)^{-\min_{r,j}\alpha_{r,j}/p}$
We now combine the two sources of error using the triangle inequality:
$\lVert f_{0}-\hat{f}\rVert_{\infty}$ $\leq\lVert f_{0}-f_{\text{ideal}}\rVert_{\infty}+\lVert f_{\text{ideal}}-\hat{f}\rVert_{\infty}$
$\leq C^{\prime}N^{-\min_{r,j}\alpha_{r,j}}+C^{\prime\prime}(D^{\star}/R)^{-\min_{r,j}\alpha_{r,j}/p}$
Substituting $N\asymp(D^{\star}/R)^{1/p}$ into the first term, we get:
$\lVert f_{0}-\hat{f}\rVert_{\infty}$ $\leq C^{\prime}\left((D^{\star}/R)^{1/p}\right)^{-\min_{r,j}\alpha_{r,j}}+C^{\prime\prime}(D^{\star}/R)^{-\min_{r,j}\alpha_{r,j}/p}$
$=C^{\prime}(D^{\star}/R)^{-\min_{r,j}\alpha_{r,j}/p}+C^{\prime\prime}(D^{\star}/R)^{-\min_{r,j}\alpha_{r,j}/p}$
$=(C^{\prime}+C^{\prime\prime})R^{\min_{r,j}\alpha_{r,j}/p}\cdot D^{\star\ -\min_{r,j}\alpha_{r,j}/p}$
Letting $C=(C^{\prime}+C^{\prime\prime})R^{\min_{r,j}\alpha_{r,j}/p}$, which is a constant with respect to $D^{\star}$, we arrive at the final result:
$\lVert f_{0}-\hat{f}\rVert_{\infty}\leq CD^{\star\ -\min_{r,j}\alpha_{r,j}/p}$
This completes the proof of the global ridge approximation lemma. ∎
@@ C.2 Proof of Theorem 1
    The split-net $\mathcal{Z}$ is $(\chi_{0}^{*},c_{n})$-dense, where $c_{n}\lesssim(\epsilon_{n}/(\lambda p))^{1/\alpha_{\min}}$.
@@
    The split net is $(\Xi_{r}^{*},\bm{\alpha}_{r},L)$-regular for every r=1,2,,Rr=1,2,...,R, with $2^{-L}\leq(\epsilon_{n}/D)^{p/\alpha_{\min}}$.
@@
  • (A8)
    For the local ridge function $f_{D}^{r}(\bm{x})$ approximating $f_{0}$ on $\Xi_{r}^{*},$ we assume $\sum_{l=1}^{D}|\beta_{l}^{(r)}|\leq C_{\text{outer}}D$ and $\|\bm{\omega}_{l}^{(r)}\|_{1}\leq C_{\text{inner}}$ for constants $C_{\text{outer}}>0$ and $C_{\text{inner}}>0$.
These assumptions ensure that the boundaries of the underlying box partition can be well detected by the binary tree partitioning rule. In particular, assumptions (A4) and (A6) ensures the split-net is dense enough to approximate the unknown partition’s boundaries. Assumption (A5) sets the anisotropic k-d tree depth $L$ so that the partition can achieve the correct approximation rate w.r.t. $\bm{\alpha}_{r}$. Assumption (A7) caps the maximum depth needed for any sub-box $\Omega_{r}^{*}$ to $\lesssim\log n$, preventing excessively large trees. It is useful to state that throughout this section, we define the class of functions defined on any tree partition $\mathcal{T}$ as $\mathcal{F}_{\mathcal{T}}$. The primary role of (A8) is to control the smoothness of the constructed ridge function approximant $\hat{f}_{0}$. Specifically, it is essential for bounding the modulus of continuity of $\hat{f}_{0}$. The constants $C_{\text{outer}}$ and $C_{\text{inner}}$ are guaranteed to exist by the construction argument elaborated in Lemma C1 and Lemma 1, since $f_{0}$ is smooth, the intermediate polynomial has bounded coefficients (DeVore and Lorentz,, 1993, Chapter 8).
Remark C1.
Another implication of assumption (A8) is that the activation function $\varphi$ is bounded over its effective domain within the model. Fixing a particular ridge leaf $d$, the argument of the activation function is $z=\hat{\bm{\omega}}_{tk}^{\top}\bm{x}+\hat{b}_{tk}$. The covariates $\bm{x}$ belong to the compact set $0,1]^{p}$. By assumption (A8), the inner weights are bounded such that $\|\hat{\bm{\omega}}_{tk}\|_{1}\leq C_{\text{inner}}$. We also assume the constructed bias terms are bounded, $|\hat{b}_{tk}|\leq C_{\text{bias}}$, as is standard. Therefore, the argument $z$ is always evaluated within a compact interval [M,M][-M,M], where $M=p\cdot C_{\text{inner}}+C_{\text{bias}}$, since
$|z|=|\hat{\bm{\omega}}_{tk}^{\top}\bm{x}+\hat{b}_{tk}|\leq\|\hat{\bm{\omega}}_{tk}\|_{\infty}\|\bm{x}\|_{1}+|\hat{b}_{tk}|\leq\|\hat{\bm{\omega}}_{tk}\|_{1}\cdot p+|\hat{b}_{tk}|\leq M.$
Assumption (P4) states that $\varphi\in C^{k}(\mathbb{R})$, which implies $\varphi$ is continuous. By the Extreme Value Theorem, a continuous function on a compact set is bounded. Thus, there exists a finite constant $C_{\varphi}=\sup_{z\in[-M,M]}|\varphi(z)|$.
@@
Assume (A1)–(A8), (P1)–(P4). Then for any $f_{0}\in\mathcal{H}^{\mathcal{A},p}_{\lambda}(\bm{\chi_{0}^{*}})$, there exists an anisotropic k-d tree partition $\hat{\mathcal{T}}$ and an associated piecewise ridge function $\hat{f}_{0}\in\mathcal{F}_{\hat{\mathcal{T}}}$ such that
$\|f_{0}-\hat{f}_{0}\|_{\infty}\lesssim\epsilon_{n}\coloneqq(\lambda p)^{\frac{p}{2\bar{\alpha}+p}}\left(\frac{R\log n}{n}\right)^{\frac{\bar{\alpha}}{2\bar{\alpha}+p}}.$
@@
Fix any $\bm{x}\in[0,1]^{p}$. Let $\Omega_{rk}^{0}$ be the leaf of the approximating partition $\hat{\mathcal{T}}$ that contains $\bm{x}$, where $\Omega_{rk}^{0}$ is a subset of some true region $\Xi_{r}^{*}$. We project $\bm{x}$ onto the closest point in the true region within that leaf, defining $\bm{x}^{*}=\arg\min_{\bm{z}\in\text{cl}(\Omega_{rk}^{0}\cap\Xi_{r}^{*})}\|\bm{x}-\bm{z}\|_{1}$.
The total error is decomposed as:
$|f_{0}(\bm{x})-\hat{f}_{0}(\bm{x})|\leq\underbrace{|f_{0}(\bm{x})-f_{0}(\bm{x}^{*})|}_{(A)}+\underbrace{|f_{0}(\bm{x}^{*})-\hat{f}_{0}(\bm{x}^{*})|}_{(B)}+\underbrace{|\hat{f}_{0}(\bm{x}^{*})-\hat{f}_{0}(\bm{x})|}_{(C)}$
We bound each term separately.
Bounding (A). This term represents the error from $f_{0}$ due to the mismatch between the true partition $\Xi_{r}^{*}$ and the approximating partition leaf $\Omega_{rk}^{0}$. The distance is bounded by $\|\bm{x}-\bm{x}^{*}\|_{\infty}\leq c_{n}$. Using the modulus of continuity of $f_{0}$, we have:
$|f_{0}(\bm{x})-f_{0}(\bm{x}^{*})|$ $\leq\lambda\sum_{j=1}^{p}|x_{j}-x_{j}^{*}|^{\alpha_{rj}}$
$\leq\lambda\sum_{j=1}^{p}c_{n}^{\alpha_{rj}}\leq\lambda pc_{n}^{\alpha_{\min}}$
By our assumption (A4), we explicitly choose $c_{n}\leq(\epsilon_{n}/(\lambda p))^{1/\alpha_{\min}}$. Substituting this in:
$|f_{0}(\bm{x})-f_{0}(\bm{x}^{*})|\leq\lambda p\left(\frac{\epsilon_{n}}{\lambda p}\right)^{\alpha_{\min}/\alpha_{\min}}=\epsilon_{n}.$
Thus, term (A) is $O(\epsilon_{n})$.
Bounding (B). This term is the error of our constructed ridge function approximant within the true region $\Xi_{r}^{*}$. On any leaf within $\Xi_{r}^{*}$, $\hat{f}_{0}$ is equal to a specific ridge function $f_{D}^{r}(\bm{x})$. Term (B) is $|f_{0}(\bm{x}^{*})-f_{D}^{r}(\bm{x}^{*})|$. This is the local approximation error from Lemma C1, which guarantees an error bound of $O(D^{-\bar{\alpha}/p})$. By our assumption (A3), we choose the number of ridge units per region to be $D\asymp\epsilon_{n}^{-p/\bar{\alpha}}$. This gives:
$|f_{0}(\bm{x}^{*})-\hat{f}_{0}(\bm{x}^{*})|\lesssim D^{-\bar{\alpha}/p}\asymp\left(\epsilon_{n}^{-p/\bar{\alpha}}\right)^{-\bar{\alpha}/p}=\epsilon_{n}.$
Thus, term (B) is $O(\epsilon_{n})$.
Bounding (C). This term captures how much the approximating function $\hat{f}_{0}$ changes between $\bm{x}$ and $\bm{x}^{*}$. On the given leaf, $\hat{f}_{0}=f_{D}^{r}$.
$|\hat{f}_{0}(\bm{x}^{*})-\hat{f}_{0}(\bm{x})|$ $=|f_{D}^{r}(\bm{x}^{*})-f_{D}^{r}(\bm{x})|$
$=\left|\sum_{l=1}^{D}\beta_{l}^{(r)}\left[\varphi(\bm{\omega}_{l}^{(r)\top}\bm{x}^{*}+b_{l}^{(r)})-\varphi(\bm{\omega}_{l}^{(r)\top}\bm{x}+b_{l}^{(r)})\right]\right|$
$\leq\sum_{l=1}^{D}|\beta_{l}^{(r)}|\left|\varphi(\bm{\omega}_{l}^{(r)\top}\bm{x}^{*}+b_{l}^{(r)})-\varphi(\bm{\omega}_{l}^{(r)\top}\bm{x}+b_{l}^{(r)})\right|$
Using assumption (P4), $\varphi$ is Lipschitz with constant $L_{\varphi}$:
$|\hat{f}_{0}(\bm{x}^{*})-\hat{f}_{0}(\bm{x})|$ $\leq\sum_{l=1}^{D}|\beta_{l}^{(r)}|\cdot L_{\varphi}\cdot|\bm{\omega}_{l}^{(r)\top}(\bm{x}^{*}-\bm{x})|$
$\leq\sum_{l=1}^{D}|\beta_{l}^{(r)}|\cdot L_{\varphi}\cdot\|\bm{\omega}_{l}^{(r)}\|_{1}\|\bm{x}^{*}-\bm{x}\|_{\infty}$ (Hölder’s Inequality)
Using our regularity assumption (A8):
$|\hat{f}_{0}(\bm{x}^{*})-\hat{f}_{0}(\bm{x})|\leq(C_{\text{outer}}D)\cdot L_{\varphi}\cdot C_{\text{inner}}\cdot\|\bm{x}^{*}-\bm{x}\|_{\infty}$
The distance $\|\bm{x}^{*}-\bm{x}\|_{\infty}$ is bounded by the size of the anisotropic k-d tree leaf, which is at most $2^{-L\alpha_{\min}/p}$. By our assumption (A5), we choose $L$ such that $2^{-L}\leq(\epsilon_{n}/D)^{p/\alpha_{\min}}$. Plugging this in:
$|\hat{f}_{0}(\bm{x}^{*})-\hat{f}_{0}(\bm{x})|\lesssim D\cdot 2^{-L\alpha_{\min}/p}\leq D\cdot\left(\left(\frac{\epsilon_{n}}{D}\right)^{p/\alpha_{\min}}\right)^{\alpha_{\min}/p}=D\cdot\frac{\epsilon_{n}}{D}=\epsilon_{n}.$
Thus, term (C) is $O(\epsilon_{n})$.
Combining the bounds for the three parts, we have:
$|f_{0}(\bm{x})-\hat{f}_{0}(\bm{x})|\leq|(A)|+|(B)|+|(C)|\lesssim\epsilon_{n}+\epsilon_{n}+\epsilon_{n}=O(\epsilon_{n}).$
Since this holds for any $\bm{x}\in[0,1]^{p}$, we conclude that $\|f_{0}-\hat{f}_{0}\|_{\infty}\lesssim\epsilon_{n}$. ∎